OpenStudy (anonymous):

A certain forest covers an area of 2800km^2. Suppose that each year this area decreases by 8%. What will the area be after 12years?Use the calculator provided and round your answer to the nearest square kilometer. WIll medal

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

The formula you'll use is \[\Large A = P(1+r)^t\] In this case, P = 2800, r = -0.08 and t = 12

Directrix (directrix):

2800 Loss of 8% per year is a retention of 92% per year 12 years 2800*(.92)^12= ? @bri97

Directrix (directrix):

See below regarding use of Google calculator

Directrix (directrix):

> round your answer to the nearest square kilometer.